klpoole1:

Two electric motors drive two elevators of equal mass in a three-story building 10 meters tall. Each elevator has a mass of 1,000 kg. The first elevator can do this work in 5.0 seconds. Calculate the power output of the first motor. 2.0 x 102 W 2.0 x 103 W 2.0 x 104 W 2.0 x 105 W

umm:

The answer is 19620 Watt = 19.62 kWatt Power output (P) is work (W) divided by time (t): P = W/t A work (W) is a product of a force (F) and a distance (d): W = F · d Since the force is a product of mass and acceleration: F = m · a then work would be: W = m · a · h P = W/t W = m · a · h P = m · a · h / t We have: P = ? m = 1000 kg a = 9.81 m/s² (gravitational acceleration on the Earth) h = 10 m t = 5s P = m · a · h / t = 1000 kg · 9.81 m/s² · 10 m / 5 s = 19620 Watt = 19.62 kWatt

umm:

\(\color{#800000}{\text{Originally Posted by: Brainly}}\)